OpenStudy (kamille):

Can anyone help me? There is an geometrical progression. The sum of first members of the sequance can be calculated according to the formula: \[S _{n}=\frac{ 1 }{ 4 }(2^{n}-1)\] Need to find a) sixth member of a sequance (\b_{6}\) and "q" So, i think I can do by this: \[b _{6}=S _{6}-S _{5}=(\frac{ 1 }{ 4 }(2^{6}-1))-((\frac{ 1 }{ 4 }(2^{5}-1))=15,75-8,25=7,5\] but the answwer should be 7, any suggestions?

OpenStudy (kamille):

@saifoo.khan

OpenStudy (kamille):

Maybe you don't understand my labelling? I know it is quite different in other countries?

OpenStudy (kamille):

@hartnn any ideas?:/

hartnn (hartnn):

S6-S5 = 1/4 {2^6-1-2^5+1} = 1/4 (32) = 8 .....

OpenStudy (kamille):

Oh, right ((14(2^5−1))=7.75 and the answer should be 8... Well, maybe mistake in the answer sheet:/

hartnn (hartnn):

even i think so...

OpenStudy (kamille):

thank you!